# Statistical Inference for Pairwise Comparison Models

### v2

@@ 7 Discussion
Despite the broad applicability of the established result in terms of model parameterization, there are practical considerations that motivate further exploration of the problem. For example, in scenarios where multiple data sources are available for the same subjects of interest, different pairwise comparison models might be employed. A promising avenue for future investigation involves integrating the estimated information derived from different yet related models. The second aspect pertains to temporal dynamics. As can be seen in the ATP example, the data span nearly fifty years. In such contexts, it is more appropriate to also account for the temporal effects on the statistical models employed to gauge comparison outcomes and on the sampling of the graph. Addressing these considerations may involve incorporating additional covariate information. We intend to pursue these directions in our future work.
Appendices
Appendix A Proof of Lemma 13
@@
Lemma 14
If $\lim_{n\to\infty}(\kappa^{7}_{n}q^{6}_{n}\log n)/(np^{7}_{n})=0$, then there exists an absolute constant $C$ such that $\|{\mathcal{A}}-\mathcal{P}_{1}\|_{2}\leq 1-C\{p_{n}/(\kappa_{n}q_{n})\}^{2}$.
Lemma 15
@@
${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (13)
@@
$-\mathcal{P}_{1}{\mathcal{D}}^{-1/2}\bar{H}({\bm{W}})({\widehat{\bm{u}}}-{\bm{u}}^{*})=\frac{1}{\sum_{i\in[n]}{\mathcal{D}}_{ii}}({\mathcal{D}}_{11}^{1/2},\ldots,{\mathcal{D}}_{nn}^{1/2})^{\top}\bm{1}^{\top}_{n}(\nabla l({\bm{u}}^{*})-\nabla l({\widehat{\bm{u}}}))$
@@
which follows as a result of the validity assumption on $f$ (that implies $g(x;y)+g(-x;-y)=0$). Combining both pieces, we obtain $\mathcal{P}_{1}\bm{\zeta}=0$. Meanwhile, since ${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1}$, it follows from the Cauchy–Schwarz inequality that
$\|{\mathcal{L}}^{\dagger}_{sym}\bm{\zeta}\|_{\infty}=$ $\left\|\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}\bm{\zeta}\right\|_{\infty}$ (14)
@@
If choosing $L=C(\kappa_{n}q_{n}/p_{n})^{2}\log n$ for some sufficiently large constant $C>0$, then the second term in (15) is $o_{p}(1)$ as a result of Lemma 14. For the first term in (15), it follows from the direct computation using (13) that ${\mathcal{A}}^{t}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$, viewed as an operator in the $\ell_{\infty}$ norm, can be bounded as
$\|{\mathcal{A}}^{t}\|_{\ell_{\infty}\to\ell_{\infty}}$ $=\max_{i\in[n]}\sum_{j\in[n]}|\{{\mathcal{A}}^{t}\}_{ij}|$ (16)
(13)${{\leq}}\max_{i\in[n]}\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\sum_{j\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-1}j}}{{{\mathcal{D}}_{ii}^{1/2}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-1}j_{t-1}}{{\mathcal{D}}_{jj}^{1/2}}}$
@@
$=\frac{\max_{i\in[n]}{{\mathcal{D}}_{ii}^{1/2}}}{\min_{k\in[n]}{{\mathcal{D}}_{kk}^{1/2}}}\lesssim\left(\frac{\kappa_{n}q_{n}}{p_{n}}\right)^{1/2}.$
Consequently,
@@
To prove the last statement in Lemma 13, note that since $({\mathcal{A}}-\mathcal{P}_{1})^{t}$ is semi-positive definite, $\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\leq\|({\mathcal{A}}-\mathcal{P}_{1})^{t}\|_{2}$ for every $t$. By a similar truncation argument as above, we only need to prove for summation of $t$ from $2$ to $L$, and this can be verified by a direct computation as in (16):
$\max_{i\in[n]}\sum_{t=2}^{L}\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\lesssim\frac{L}{({c_{n,3}np_{n}})^{1/2}}\lesssim\frac{c_{n,4}^{2}}{c_{n,3}^{5/2}}\left\{\frac{q_{n}^{4}(\log n)^{2}}{np_{n}^{5}}\right\}^{1/2}=o_{p}(1),$
where the details are omitted.
Appendix B Proofs of Lemmas 1415
Before showing Lemmas 1415, we first cite the uniform consistency result in Han et al. (2023):
Lemma 16
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 4-6, if
@@ B.1 Proof of Lemma 14
Denote eigenvalues of $\mathcal{L}_{\text{sym }}=I-\mathcal{A}=I-\mathcal{D}^{-1/2}\mathcal{W}\mathcal{D}^{-1/2}$ in increasing order as $0=\lambda_{1}\leq\cdots\leq\lambda_{n}\leq 2$, where ${\mathcal{D}}$ and $\mathcal{W}$ are all random matrices (where randomness is from the comparison graph ${\mathcal{G}}$ only). In particular, $\mathcal{W}$ is an off-diagonal weight matrix with $\mathcal{W}_{ij}=\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}z_{ij}$, where
@@
due to Assumption 6, with ${\mathbb{E}}_{X_{ij}}[\cdot]$ denoting expectation for $X_{ij}$.
It is clear $\left\|\mathcal{A}-\mathcal{P}_{1}\right\|_{2}=\max\left\{1-\lambda_{2},\lambda_{n}-1\right\}$.
To analyze the spectrum of $\mathcal{L}_{\text{sym }}$, we consider its expected version first. Let $\overline{\mathcal{D}}$ and $\overline{\mathcal{W}}$ be the expectation of $\mathcal{D}$ and $\mathcal{W}$, respectively, and ¯$\bar{\mathcal{L}}_{\text{sym }}=I-\bar{\mathcal{D}}^{-1/2}\bar{\mathcal{W}}\bar{\mathcal{D}}^{-1/2}$. By definition, ¯$\bar{\mathcal{L}}_{\text{sym }}$ is also a normalized graph Laplacian matrix; we denote the eigenvalues of ¯$\overline{\mathcal{L}}_{\text{sym }}$ by $0=\bar{\lambda}_{1}\leq\cdots\leq\bar{\lambda}_{n}\leq 2$. A computation based on Cheeger’s inequalities yields a lower bound for the spectral gap of the expected Laplacian (Chung, 1997; Bauer and Jost, 2013; Han and Xu, 2023): There exists an absolute constant $c>0$ such that
$\max\left\{1-\bar{\lambda}_{2},\bar{\lambda}_{n}-1\right\}\leq 1-c\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2}.$ (17)
We next apply a concentration argument to show that the spectra of $\mathcal{L}_{\text{sym }}$ and ¯$\bar{\mathcal{L}}_{\text{sym }}$ are close. To this end, we first apply Weyl’s inequality to obtain
$\max\left\{|(1-\lambda_{2})-(1-\bar{\lambda}_{2})|,|(\lambda_{n}-1)-(\bar{\lambda}_{n}-1)|\right\}$
$\leq$ $\left\|\mathcal{L}_{\text{sym }}-\overline{\mathcal{L}}_{\text{sym }}\right\|_{2}=\left\|\mathcal{D}^{-1}\mathcal{W}-\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}$
$\leq$ $\left\|\left(\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right)\overline{\mathcal{W}}\right\|_{2}+\left\|\mathcal{D}^{-1}(\mathcal{W}-\overline{\mathcal{W}})\right\|_{2}$
$\leq$ $\left\|\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right\|_{2}\|\overline{\mathcal{W}}\|_{2}+\left\|\mathcal{D}^{-1}\right\|_{2}\|\mathcal{W}-\overline{\mathcal{W}}\|_{2}.$
@@
$\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\lesssim\left({c_{n,4}\overline{\mathcal{D}}_{ii}\log n}\right)^{1/2}\quad i\in[n],$
@@
$\left\|\mathcal{D}^{-1}-\overline{\mathcal{D}}^{-1}\right\|_{2}$ $=\max_{i\in[n]}\left|\mathcal{D}_{ii}^{-1}-\overline{\mathcal{D}}_{ii}^{-1}\right|\lesssim\max_{i\in[n]}\left(\overline{\mathcal{D}}_{ii}^{-2}\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\right)\lesssim\left\{\frac{c_{n,4}\log n}{(c_{n,3}np_{n})^{3}}\right\}^{1/2}$
$\left\|\mathcal{D}^{-1}\right\|_{2}$ $\lesssim\left\|\overline{\mathcal{D}}^{-1}\right\|_{2}\lesssim\frac{1}{c_{n,3}np_{n}}.$
@@
$\|\overline{\mathcal{W}}\|_{2}=\left\|\overline{\mathcal{D}}\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}\leq\|\overline{\mathcal{D}}\|_{2}\left\|\overline{\mathcal{D}}^{-1}\overline{\mathcal{W}}\right\|_{2}\leq 2\|\overline{\mathcal{D}}\|_{2}=2\max_{i\in[n]}\overline{\mathcal{D}}_{ii}\lesssim c_{n,4}nq_{n}.$
By the matrix Bernstein inequality (Tropp, 2012), with probability at least $1-n^{-2}$,
@@
$\left\|\mathcal{L}_{\text{sym }}-\overline{\mathcal{L}}_{\text{sym}}\right\|_{2}\lesssim\left(\frac{\kappa_{n}^{3}q_{n}^{2}\log n}{np_{n}^{3}}\right)^{1/2}+\left(\frac{\kappa_{n}^{2}q_{n}\log n}{np_{n}^{2}}\right)^{1/2}.$ (18)
The desired result follows if the expected spectral gap in (17) dominates the concentration deviations (18):
@@ B.2 Proof of Lemma 15
$\{H({\bm{u}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})=-\partial_{2}g(X_{ij},u_{i}-u_{j})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j\end{array}\right.$ (21)
$\{\bar{H}({\bm{W}})\}_{ij}$ =$=\left\{\begin{array}[]{ll}\partial_{ij}l(\bm{w}_{i})=-\partial_{2}g(X_{ij},w_{ii}-w_{ij})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l(\bm{w}_{i})&\quad i=j\end{array}\right.$ (24)
where ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ points on the line segment between ${\widehat{\bm{u}}}$ and ${\bm{u}}^{*}$, and ${\bm{w}}_{i}=(w_{i1},\ldots,w_{in})\in{\mathbb{R}}^{n}$. Write the $i$th element of $\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)$ as
|[$\left|\left[\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|$
@@
|$\left|\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}_{ij}\right|$ =(21)|{${{=}}\left|\left\{\partial_{2}g(X_{ij},u^{*}_{i}-u^{*}_{j})-\partial_{2}g(X_{ij},w_{ii}-w_{ij})\right\}\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}\right|$
@@
|[$\left|\left[\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|\leq 4c_{n,5}\|\widehat{{\bm{u}}}-{\bm{u}}^{*}\|^{2}_{\infty}|\delta\{i\}|.$
@@
The proof of the first statement is completed. Next, we prove the second statement in Lemma 15. Similar to the previous proof, we can decompose |[$\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$ into two parts and obtain
|[$\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$
@@
[…]
@@
$(A.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}q_{n}(\log n)^{2}}{c_{n,3}np_{n}^{3/2}}.$ (25)
[…]
$\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}$ and $\widehat{u}_{j}-u^{*}_{j}$. To address this issue, we apply a leave-one-out approach based on a modification of Gao et al. (2023). For $i\in[n]$, let
$l^{(-i)}({\bm{u}})=\sum_{j\neq i,k\neq i,(j,k)\in{\mathcal{E}}}\log f(X_{jk},u_{j}-u_{k}).$ (26)
@@
[…]
$(J.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}.$ (28)
[…]
$\lesssim c_{n,4}\left({|\delta\{i\}|}\right)^{1/2}\times\|{\widehat{\bm{u}}}_{-i}-{\widehat{\bm{u}}}^{(-i)}\|_{2},$ (29)
@@
$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\leq\frac{\lVert\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\lVert_{2}}{\lambda_{2}\{-\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}},$ (30)

### v3

@@ 7 Discussion
Despite the broad applicability of the established result in terms of model parameterization, there are practical considerations that motivate further exploration of the problem. For example, in scenarios where multiple data sources are available for the same subjects of interest, different pairwise comparison models might be employed. A promising avenue for future investigation involves integrating the estimated information derived from different yet related models. The second aspect pertains to temporal dynamics. As can be seen in the ATP example, the data span nearly fifty years. In such contexts, it is more appropriate to also account for the temporal effects on the statistical models employed to gauge comparison outcomes and on the sampling of the graph. Addressing these considerations may involve incorporating additional covariate information.
In addition, one theoretical research direction is to apply our technique for establishing asymptotic normality, that is, the representation of the Fisher information matrix as a weighted graph Laplacian, to a broader class of network models. In particular, we plan to adapt it to the $\beta$-model (chatterjee2011random), which shares structural similarities with the BT model, and its generalized variants in econometric networks (graham2017econometric; chen2021analysis). These investigations will form part of our future work.
Appendix A Appendix: Proof of Lemma 1
@@
Lemma 2.
If $\lim_{n\to\infty}(\kappa^{7}_{n}q^{6}_{n}\log n)/(np^{7}_{n})=0$, then there exists an absolute constant $C>0$ such that $\|{\mathcal{A}}-\mathcal{P}_{1}\|_{2}\leq 1-C\{p_{n}/(\kappa_{n}q_{n})\}^{2}$ holds with probability tending to one.
Lemma 3.
@@
${\mathcal{D}}_{ii}=\sum_{j\in\delta\{i\}}H_{ij},\quad{\mathcal{A}}_{ij}=\frac{H_{ij}}{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}},\quad\{\mathcal{P}_{1}\}_{ij}=\frac{{{\mathcal{D}}_{ii}^{1/2}{\mathcal{D}}_{jj}^{1/2}}}{\sum_{k\in[n]}{\mathcal{D}}_{kk}}.$ (A.1)
@@
$)\displaystyle-\mathcal{P}_{1}{\mathcal{D}}^{-1/2}\bar{H}({\bm{W}})({\widehat{\bm{u}}}-{\bm{u}}^{*})=\frac{1}{\sum_{i\in[n]}{\mathcal{D}}_{ii}}({\mathcal{D}}_{11}^{1/2},\ldots,{\mathcal{D}}_{nn}^{1/2})^{\top}\bm{1}^{\top}_{n}(\nabla l({\bm{u}}^{*})-\nabla l({\widehat{\bm{u}}}$))
@@
which follows as a result of the validity assumption on $f$ (that implies g(x;y)+g(x;y)=0g(x;y)+g(-x;-y)=0). Combining both pieces, we obtain $\mathcal{P}_{1}\bm{\zeta}=0$. Meanwhile, since ${\mathcal{L}}_{sym}^{\dagger}=\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}-\mathcal{P}_{1}$, it follows from the Cauchy–Schwarz inequality that
$\|{\mathcal{L}}^{\dagger}_{sym}\bm{\zeta}\|_{\infty}=\$$\left\|\sum_{t=0}^{\infty}({\mathcal{A}}-\mathcal{P}_{1})^{t}\bm{\zeta}\right\|_{\infty}$ (A.2)
@@
If choosing $L=C(\kappa_{n}q_{n}/p_{n})^{2}\log n$ for some sufficiently large constant $C>0$, then the second term in (A.3) is $o_{p}(1)$ as a result of Lemma 2. For the first term in (A.3), it follows from the direct computation using (A.1) that ${\mathcal{A}}^{t}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$, viewed as an operator in the $\ell_{\infty}$ norm, can be bounded as
$\|{\mathcal{A}}^{t}\|_{\ell_{\infty}\to\ell_{\infty}}$ $=\max_{i\in[n]}\sum_{j\in[n]}|\{{\mathcal{A}}^{t}\}_{ij}|$ (A.4)
[…]
@@
$=\frac{\max_{i\in[n]}{{\mathcal{D}}_{ii}^{1/2}}}{\min_{k\in[n]}{{\mathcal{D}}_{kk}^{1/2}}}\lesssim\left(\frac{\kappa_{n}q_{n}}{p_{n}}\right)^{1/2},$
where the last inequality follows from the standard degree concentration bounds (B.6). Consequently,
@@
To prove the last statement in Lemma 1, note that $\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\leq\|({\mathcal{A}}-\mathcal{P}_{1})^{t}\|_{2}$ for every $t$. By a similar truncation argument as above, we only need to prove the asymptotic negligibility of the partial sum of $t$ from $2$ to $L$:
$\max_{i\in[n]}\sum_{t=2}^{L}\{({\mathcal{A}}-\mathcal{P}_{1})^{t}\}_{ii}\leq\max_{i\in[n]}\sum_{t=2}^{L}\{{\mathcal{A}}^{t}\}_{ii}\lesssim\frac{L\kappa_{n}^{1/2}}{(np_{n})^{1/2}}\lesssim{\kappa_{n}^{5/2}}\left\{\frac{q_{n}^{4}(\log n)^{2}}{np_{n}^{5}}\right\}^{1/2}=o_{p}(1),$
where the first inequality follows from the fact that $({\mathcal{A}}-\mathcal{P}_{1})^{t}\preceq{\mathcal{A}}^{t}$ for $t\geq 1$, and the second inequality follows from similar computation as in (A.4), that is,
$\{{\mathcal{A}}^{t}\}_{ii}$ ${\leq}\max_{i\in[n]}\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-1}i}}{{{\mathcal{D}}_{ii}^{1/2}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-1}j_{t-1}}{{\mathcal{D}}_{ii}^{1/2}}}$
$\leq\max_{i\in[n]}(\max_{j_{t-1}\in[n]}\frac{H_{j_{t-1}i}}{{{\mathcal{D}}_{j_{t-1}j_{t-1}}}})\sum_{j_{1}\in[n]}\cdots\sum_{j_{t-1}\in[n]}\frac{H_{ij_{1}}\cdots H_{j_{t-2}j_{t-1}}}{{{\mathcal{D}}_{ii}}{\mathcal{D}}_{j_{1}j_{1}}\cdots{\mathcal{D}}_{j_{t-2}j_{t-2}}}$
[…]
Appendix B Proofs of Lemmas 2-3
Before showing Lemmas 2-3, we first cite the uniform consistency result in han2022general:
Lemma 4.
Let ${\mathcal{G}}\sim{\mathsf{G}}(n,p_{n},q_{n})$. Under Assumptions 1-3, if
@@ B.1 Proof of Lemma 2
The proof of Lemma 2 follows a similar approach as (han2023unified, Lemma 5.3). Denote the eigenvalues of $\mathcal{L}_{\text{sym }}=I-\mathcal{A}=I-\mathcal{D}^{-1/2}\mathcal{W}\mathcal{D}^{-1/2}$ in increasing order as $0=\lambda_{1}\leq\cdots\leq\lambda_{n}\leq 2$, where $\mathcal{D}$} and $\mathcal{W}$ are all random matrices (where randomness is from the comparison graph $\mathcal{G}$} only). In particular, $\mathcal{W}$ is an off-diagonal weight matrix with $\mathcal{W}_{ij}=\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}z_{ij}$, where
@@
due to Assumption 3, with ${\mathbb{E}}_{X_{ij}}[\cdot]$ denoting expectation for $X_{ij}$. It is clear $\left\|\mathcal{A}-\mathcal{P}_{1}\right\|_{2}=\max\left\{1-\lambda_{2},\lambda_{n}-1\right\}$.
[…]
$\max\left\{1-\bar{\lambda}_{2},\bar{\lambda}_{n}-1\right\}\leq 1-c\left(\frac{p_{n}}{\kappa_{n}q_{n}}\right)^{2}.$ (B.5)
[…]
$\max\left\{\left|(1-\lambda_{2})-(1-\bar{\lambda}_{2})\right|,\ \left|(\lambda_{n}-1)-(\bar{\lambda}_{n}-1)\right|\right\}$
[…]
$+\left\|\overline{\mathcal{D}}^{-1/2}\right\|_{2}\left\|\overline{\mathcal{W}}\right\|_{2}\left\|\mathcal{D}^{-1/2}-\overline{\mathcal{D}}^{-1/2}\right\|_{2}$
@@
$\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\lesssim\left({c_{n,4}\overline{\mathcal{D}}_{ii}\log n}\right)^{1/2}\quad i\in[n],$ (B.6)
@@
$\left\|\mathcal{D}^{-1/2}-\overline{\mathcal{D}}^{-1/2}\right\|_{2}$ $=\max_{i\in[n]}\left|\mathcal{D}_{ii}^{-1/2}-\overline{\mathcal{D}}_{ii}^{-1/2}\right|\lesssim\max_{i\in[n]}\left(\overline{\mathcal{D}}_{ii}^{-3/2}\left|\mathcal{D}_{ii}-\overline{\mathcal{D}}_{ii}\right|\right)\lesssim\left\{\frac{c_{n,4}\log n}{(c_{n,3}np_{n})^{2}}\right\}^{1/2}$
$\left\|\mathcal{D}^{-1/2}\right\|_{2}$ $\lesssim\left\|\overline{\mathcal{D}}^{-1/2}\right\|_{2}\lesssim\sqrt{\frac{1}{c_{n,3}np_{n}}}.$
@@
$\|\overline{\mathcal{W}}\|_{2}=\left\|\overline{\mathcal{D}}^{1/2}(I-\bar{\mathcal{L}}_{\text{sym}})\overline{\mathcal{D}}^{1/2}\right\|_{2}\leq\|\overline{\mathcal{D}}\|_{2}\left\|I-\bar{\mathcal{L}}_{\text{sym}}\right\|_{2}\leq\|\overline{\mathcal{D}}\|_{2}=\max_{i\in[n]}\overline{\mathcal{D}}_{ii}\lesssim c_{n,4}nq_{n}.$
By the matrix Bernstein inequality (tropp2012user), with probability at least $1-n^{-2}$,
@@
$\left\|\mathcal{L}_{\text{sym }}-\overline{\mathcal{L}}_{\text{sym}}\right\|_{2}\lesssim\left(\frac{\kappa_{n}^{3}q_{n}^{2}\log n}{np_{n}^{3}}\right)^{1/2}+\left(\frac{\kappa_{n}^{2}q_{n}\log n}{np_{n}^{2}}\right)^{1/2}.$ (B.7)
The desired result follows if the expected spectral gap in (B.5) dominates the concentration deviations (B.7):
@@ B.2 Proof of Lemma 3
$\{H({\bm{u}})\}_{ij}$ $=\left\{\begin{array}[]{ll}\partial_{ij}l({\bm{u}})=-\partial_{2}g(X_{ij},u_{i}-u_{j})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l({\bm{u}})&\quad i=j\end{array}\right$. (B.10)
$\{\bar{H}({\bm{W}})\}_{ij}$ $=\left\{\begin{array}[]{ll}\partial_{ij}l(\bm{w}_{i})=-\partial_{2}g(X_{ij},w_{ii}-w_{ij})\times\mathbb{I}_{\{(i,j)\in{\mathcal{E}}\}}&\quad i\neq j\\ -\sum_{k\in\delta\{i\}}\partial_{ik}l(\bm{w}_{i})&\quad i=j\end{array}\right$. (B.13)
where ${\bm{w}}_{1},\ldots,{\bm{w}}_{n}$ are $n$ points on the line segment between $^{\widehat{\bm{u}}}$ and $\bm{u}$$^{*}$, and ${\bm{w}}_{i}=(w_{i1},\ldots,w_{in})\in{\mathbb{R}}^{n}$. Write the $i$th element of $\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)$ as
$\left|\left[\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|$
@@
[…]
@@
$\left|\left[\left\{\bar{H}({\bm{W}})-H\left({\bm{u}}^{*}\right)\right\}\left(\widehat{{\bm{u}}}-{\bm{u}}^{*}\right)\right]_{i}\right|\leq 4c_{n,5}\|\widehat{{\bm{u}}}-{\bm{u}}^{*}\|^{2}_{\infty}|\delta\{i\}|.$
@@
The proof of the first statement is completed. Next, we prove the second statement in Lemma 3. Similar to the previous proof, we can decompose $\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$ into two parts and obtain
$\left\lvert[\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\}({\widehat{\bm{u}}}-{\bm{u}}^{*})]_{i}\right\lvert$
@@
To bound $(A.2)$, note $\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}$ is a sum of independent centered random variables bounded by $2c_{n,4}$. By Hoeffding’s inequality, with probability at least $1-n^{-2}$,
@@
$(A.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}-{\bm{u}}^{\ast}\|_{\infty}\lesssim\left({|\delta\{i\}|}\right)^{1/2}\times\frac{c_{n,2}c_{n,4}q_{n}(\log n)^{2}}{c_{n,3}np_{n}^{3/2}}.$ (B.14)
Unfortunately, $(A.1)$ cannot be treated similarly as a result of dependence between
$\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}$ and $\widehat{u}_{j}-u^{*}_{j}$. To address this issue, we apply a leave-one-out approach based on a modification of gao2023uncertainty. For $i\in[n]$, let
$l^{(-i)}({\bm{u}})=\sum_{j\neq i,k\neq i,(j,k)\in{\mathcal{E}}}\log f(X_{jk},u_{j}-u_{k}).$ (B.15)
@@
Now the term $(A.1)$ can be bounded using the triangle inequality as
$(A.1)\leq$ $\underbrace{\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}\left(\widehat{u}_{j}-\widehat{u}_{j}^{(i)}\right)\right|}_{(J.1)}+\underbrace{\left|\sum_{j\neq i}\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{ij}\left(\widehat{u}^{(i)}_{j}-u_{j}^{\ast}\right)\right|}_{(J.2)}.$ (B.16)
Since ${\widehat{\bm{u}}}^{(-i)}$ is independent of $\{X_{ij}\}_{j\in\delta\{i\}}$, we can bound ($J.2$) similar to the steps in bounding $(A.2)$ by first conditioning on ${\widehat{\bm{u}}}^{(-i)}$. With probability at least $1-n^{-2}$,
$(J.2)\lesssim c_{n,4}\left(|\delta\{i\}|\log n\right)^{1/2}\times\|{\widehat{\bm{u}}}^{(-i)}-{\bm{u}}_{-i}^{\ast}\|_{\infty}.$ (B.17)
To bound ($J.1$), we apply the Cauchy–Schwarz inequality to obtain
$(J.1)$ $\leq\left\|\left[\left\{H({\bm{u}}^{*})-H^{\ast}({\bm{u}}^{\ast})\right\}_{i,:}\right]_{-i}\right\|_{2}\|{\widehat{\bm{u}}}_{-i}-{\widehat{\bm{u}}}^{(-i)}\|_{2}$
$\lesssim c_{n,4}\left({|\delta\{i\}|}\right)^{1/2}\times\|{\widehat{\bm{u}}}_{-i}-{\widehat{\bm{u}}}^{(-i)}\|_{2},$ (B.18)
@@
$\|\widehat{{\bm{u}}}^{(-i)}-\widehat{{\bm{u}}}_{-i}\|_{2}\leq\frac{\lVert\nabla l^{(-i)}(\widehat{{\bm{u}}}_{-i})\lVert_{2}}{\lambda_{2}\{-\nabla^{2}l^{(-i)}(\widetilde{{\bm{u}}})\}}$, (B.19)